# Towards Reliable Forgetting: A Survey on Machine Unlearning Verification, Challenges, and Future Directions

### v1

@@ 5.3.2. Reproducible Unlearning Verification
Q1: Theoretical Guarantee & Q2: Access Requirement These methods offer strong theoretical guarantees by ensuring that model updates and checkpoints can be validated against expected changes, making the verification process theoretically sound (Q1: ). However, they typically require access to internal model parameters, weights, and training logs—resources often unavailable in commercial settings—thus limiting their practical applicability (Q2: ).
Q3: Sample-Level Verification & Q4: Verification Accuracy. Reproducible methods support fine-grained verification at the individual sample level, ensuring that forgotten data is excluded during retraining (Q3: ). PoUL achieves high verification accuracy through its repeatability and precise tracking of model updates (Q4: ). Cryptographic extensions, such as those using Trusted Execution Environments (TEEs) or SNARKs, further enhance trust by providing tamper-proof proofs of unlearning (Q4: ).
Q5: No Pre-Injected Data Required, Q6: Efficiency and Scalability, & Q7: Method Specificity. These methods do not rely on specially crafted or pre-injected data; as long as relevant training and update records are maintained, they can be applied across various unlearning scenarios with flexibility (Q5: ). Nonetheless, reproducing model checkpoints and verifying updates can be computationally expensive, especially for large models or datasets, which limits scalability (Q6: ). Moreover, while the PoUL method exhibits broad applicability across various unlearning algorithms (Q7: ), its cryptographic variants, despite enhancing verification reliability, may reduce generality. For instance, the approach proposed in (Weng et al., 2024) is primarily limited to SISA-based settings (Q7: ).
Table 7. Comparison of unlearning verification methods.The table presents the evaluation of fingerprint verification methods across seven key dimensions: Q1-Q7. Each method category is scored using the following symbols:  (fully satisfies),  (does not satisfy),  (mostly satisfies, with some caveats),  (Others). The seven dimensions evaluated are as follows: Q1: Theoretical Guarantee, Q2: Access Requirement, Q3: Sample-Level Verification, Q4: Verification Accuracy, Q5: No Pre-Injected Data Required, Q6: Efficiency and Scalability, and Q7: Method Specificity.
@@ 5.4. Comparison with Different Verification
Behavioral methods are generally effective in practical deployment due to their low access requirements and high computational efficiency, making them particularly suitable for black-box environments such as commercial APIs. However, they lack formal guarantees, offer limited support for sample-level verification, and often exhibit instability in real-world scenarios. For instance, fingerprint techniques, such as adversarial fingerprints or backdoor watermarking, may fail to reliably activate the embedded signals when model architectures or data distributions shift, resulting in false positives or negatives. Privacy analysis methods also vary significantly in performance across different models and tasks, with inference accuracy often degrading after model updates. Similarly, performance methods, including relearn time, are sensitive to factors such as random initialization and hyperparameter tuning, leading to inconsistent results and reduced reproducibility.
@@
6. Threats to Verification Reliability.
@@ 6.2. Threats from the Parameter Space
$\left\|g(w_{i},d^{t})-g(w_{i},\tilde{d}^{(t)})\right\|\leq\varepsilon.$
@@
To further improve attack efficiency, Zhang et al. (Zhang et al., 2024a) propose an input-space nearest neighbor substitution strategy. Rather than minimizing the gradient deviation, this approach directly replaces each forgotten sample in a minibatch with the most similar retained sample from the same class, based on Euclidean distance. For each forgotten point $(x,y)\in D_{f}$, the attacker selects:
$N(x,y)=\arg\min_{(x^{*},y^{*})\in D\setminus D_{f},\,y^{*}=y}\|x^{*}-x\|.$
@@
To mitigate these risks, some recent works (Eisenhofer et al., 2022; Weng et al., 2024) propose incorporating cryptographic enhancements into the PoUL framework, such as trusted execution environments or zero-knowledge proofs. While these techniques improve integrity and tamper-resistance, their significant computational and deployment overhead limits their practicality in large-scale or resource-constrained settings.
@@ 6.3. Threats from the Behavioral Space
Suppressing Behavioral Fingerprints.
Verification methods based on fingerprint detection rely on the model’s distinctive responses to crafted probes to determine whether remnants of forgotten data persist. However, such behavioral signatures can be neutralized. Adversaries may train the model adversarially to suppress fingerprint activation (Bai et al., 2021; Jia et al., 2022), or apply input filtering and sanitization mechanisms to block suspicious fingerprint queries (Wang et al., 2022; Abusnaina et al., 2021). Targeted fingerprint removal techniques (Zheng et al., 2022; Guo et al., 2020) can eliminate verification triggers while preserving memorized content, thereby invalidating the behavioral cues relied upon by the auditor.
Obfuscating Privacy Leakage Signals.
Privacy analysis verification relies on the success of inference attacks to infer whether forgotten data still influences the model. However, these signals can be intentionally obfuscated. For instance, retraining with differential privacy (Jia et al., 2019), injecting noise into gradients, or clipping model outputs (Gong et al., 2023; Wen et al., 2021) can reduce attack effectiveness. Consequently, models may appear to have forgotten certain data while covertly retaining sensitive information in internal representations or decision boundaries.erhead limits their scalability in real-world settings.
Spoofing Performance Degradation.
Model performance methods monitor metrics such as accuracy decline, relearn time, or loss recovery to infer whether unlearning has occurred. Yet these indicators can be spoofed by adversaries. Through data poisoning (Tolpegin et al., 2020; Fang et al., 2020), the model’s performance can be manipulated to mimic the expected degradation or improvement patterns. In such cases, the model passes the performance check without performing actual data removal, undermining the verification’s trustworthiness.
@@ 7. Open Questions and Future Directions
A fundamental challenge in unlearning research is the absence of a unified and operational definition of forgetting. Current approaches rely on diverse criteria, including fingerprint (Zhang et al., 2024d; Tam et al., 2024; Goel et al., 2022), privacy analysis (Tu et al., 2024; Di et al., 2024; Sharma et al., 2024), and model performance (Graves et al., 2021; Mehta et al., 2022). This definitional fragmentation creates a critical ambiguity: if a model passes certain verification protocols but fails others, can it still be considered to have truly forgotten the data? Without a common standard, service providers may selectively adopt favorable verification strategies and claim compliance without achieving the intended privacy guarantees. The lack of formal consensus not only impedes fair comparison across methods but also opens the door to unverifiable or misleading claims. Future research should aim to establish a rigorous, interpretable, and widely applicable definition of forgetting that accommodates various verification paradigms and practical deployment needs.
Reliable Unlearning Verification Methods.
Beyond definitional ambiguity, the reliability of existing verification methods remains a pressing concern. Many empirical approaches, such as those based on privacy inference or performance degradation, depend heavily on model outputs and the effectiveness of auxiliary attack modules. When these modules are poorly calibrated or overly sensitive to distributional shifts, the resulting verification can become unstable or misleading. At the same time, theoretically grounded methods, including differential privacy verification and reproducibility verification, also suffer from fundamental limitations. For instance, reproducibility verification assumes that if a model trained with and without a specific data point yields identical parameters, then the data has been successfully forgotten. However, as shown in (Thudi et al., 2022), it is possible to construct datasets that differ by a single point $x$ yet result in nearly indistinguishable models. This allows for false claims of forgetting without any actual data removal. Although recent work has explored the integration of cryptographic primitives into reproducibility verification to enhance its trustworthiness, these techniques remain computationally intensive and difficult to scale. This underscores the need for verification methods that combine theoretical rigor with practical feasibility. Such methods should provide reliable forgetting guarantees without imposing excessive computational or access requirements.
@@
Despite growing interest in unlearning verification, many existing methods are developed under overly idealized assumptions. It is often presumed that data distributions are stationary (Chen et al., 2021; Fan et al., 2023), model internals are fully accessible (Guo et al., 2019; Zhang et al., 2024b), and the training process is entirely reproducible (Thudi et al., 2022; Eisenhofer et al., 2022). These assumptions rarely hold in real-world deployments, where models are subject to stochastic training dynamics, evolving input distributions, and limited access due to privacy policies or API restrictions. Such constraints introduce substantial uncertainty that current methods are ill-prepared to handle. For instance, small perturbations in initialization or mini-batch order can significantly affect model outputs, undermining the reliability of behavior signals. Similarly, black-box environments make gradient-based or log-based verification infeasible. This gap between theoretical design and deployment reality remains one of the least addressed yet most impactful challenges. Without explicitly accounting for these constraints, verification results may become unstable, misleading, or fundamentally inapplicable. Future research must incorporate robustness into the core design of verification protocols, including tolerance to randomness, adaptability to shifting distributions, and operability under partial or noisy access. Otherwise, verification will remain a fragile component that functions only in controlled settings and silently fails in practical use.
@@
Another insufficiently explored limitation is the inconsistency of verification effectiveness across different data samples. Many existing methods, particularly those based on membership inference, perform exceptionally well on atypical or poorly generalized samples, which are the most likely to be memorized by the model (Zhong et al., 2024; Feldman and Zhang, 2020). However, they often underestimate the forgetting failure rate on well-generalized samples, leading to incomplete or biased assessments. This inconsistency reflects a deeper issue: current verification techniques often implicitly assume that all samples are equally verifiable, ignoring the heterogeneous learning dynamics inherent to modern models. Without accounting for sample-specific factors such as gradient influence, model sensitivity, or generalization difficulty, these methods risk producing misleading evaluations of unlearning effectiveness. Future work should aim to develop verification strategies that support consistent and fair assessments across a broad spectrum of data instances.
@@ 8. Conclusion
Machine unlearning is a process that enables machine learning models to forget specific training data through specialized techniques. While these techniques provide algorithm-level forgetting or approximate forgetting, the honesty of the service providers offering unlearning APIs often remains questionable. Therefore, verification becomes a crucial element to enhance the reliability of machine unlearning. However, existing work lacks a comprehensive and thorough survey of this field. To fill this gap, we present the first survey on machine unlearning verification. We classify existing verification methods into empirical and formal types, analyzing their respective advantages and limitations. Additionally, we examine the associated vulnerabilities and threats and identify several open research questions. We believe that our work can advance the study of the reliability and security of machine unlearning, thus enabling machine learning systems to more reliably handle the forgetting of sensitive data.

### v2

@@ 5.3.2. Reproducible Unlearning Verification
Q1 Theoretical Guarantee & Q2 Access Requirement. These methods offer strong theoretical guarantees by ensuring that model updates and checkpoints can be validated against expected changes, making the verification process theoretically sound (Q1. ). However, they typically require access to internal model parameters, weights, and training logs—resources often unavailable in commercial settings—thus limiting their practical applicability (Q2. ).
Q3 Sample-Level Verification & Q4 Verification Accuracy. Reproducible methods support fine-grained verification at the individual sample level, ensuring that forgotten data is excluded during retraining (Q3. ). PoUL achieves high verification accuracy through its repeatability and precise tracking of model updates (Q4. ). Cryptographic extensions, such as those using Trusted Execution Environments (TEEs) or SNARKs, further enhance trust by providing tamper-proof proofs of unlearning (Q4. ).
Q5 No Pre-Injected Data Required, Q6 Efficiency and Scalability, & Q7 Method Specificity. These methods do not rely on specially crafted or pre-injected data; as long as relevant training and update records are maintained, they can be applied across various unlearning scenarios with flexibility (Q5. ). Nonetheless, reproducing model checkpoints and verifying updates can be computationally expensive, especially for large models or datasets, which limits scalability (Q6. ). Moreover, while the PoUL method exhibits broad applicability across various unlearning algorithms (Q7. ), its cryptographic variants, despite enhancing verification reliability, may reduce generality. For instance, the approach proposed in (Weng et al., 2024) is primarily limited to SISA-based settings (Q7. ).
Table 7. Comparison of unlearning verification methods.The table presents the evaluation of verification methods across seven key dimensions: Q1-Q7. Each method category is scored using the following symbols:  (fully satisfy),  (mostly satisfies, with some caveats),  (Others),  (does not satisfy). The seven dimensions evaluated are as follows: Q1. Theoretical Guarantee, Q2. Access Requirement, Q3. Sample-Level Verification, Q4. Verification Accuracy, Q5. No Pre-Injected Data Required, Q6. Efficiency and Scalability, and Q7. Method Specificity.
@@ 5.4. Comparison with Different Verification
Behavioral methods are generally effective in practical deployment due to their low access requirements and high computational efficiency, making them particularly suitable for black-box environments such as commercial APIs. However, they lack formal guarantees, offer limited support for sample-level verification, and often exhibit instability in real-world scenarios. For instance, fingerprint techniques may fail to reliably activate the embedded signals when model architectures or data distributions shift, resulting in false positives or negatives. Privacy analysis methods also vary significantly across different models and tasks, with inference accuracy often degrading after model updates. Similarly, performance methods are sensitive to factors such as random initialization and hyperparameter tuning, leading to inconsistent results and reduced reproducibility.
@@
6. Threats to Verification Reliability
@@ 6.2. Threats from the Parameter Space
$\left\|g(\mathbf{w}_{i},d^{t})-g(\mathbf{w}_{i},\tilde{d}^{(t)})\right\|\leq\varepsilon.$
@@
To further improve attack efficiency, Zhang et al. (Zhang et al., 2024a) proposed an input-space nearest neighbor substitution strategy. Rather than minimizing the gradient deviation, this approach directly replaces each forgotten sample in a minibatch with the most similar retained sample from the same class, based on Euclidean distance. For each forgotten point $(x,y)\in D_{f}$, the attacker selects:
$N(\mathbf{x},y)=\arg\min_{(\mathbf{x}^{*},y^{*})\in D\setminus D_{f},\,y^{*}=y}\|\mathbf{x}^{*}-\mathbf{x}\|.$
@@
To mitigate these risks, some recent works (Eisenhofer et al., 2025; Weng et al., 2024) proposed incorporating cryptographic enhancements into the PoUL framework, such as trusted execution environments or zero-knowledge proofs. While these techniques improve integrity and tamper-resistance, their significant computational and deployment overhead limits their practicality in large-scale or resource-constrained settings.
@@ 6.3. Threats from the Behavioral Space
  1. (1)
    Suppressing Behavioral Fingerprints. Verification methods based on fingerprint detection rely on the model’s distinctive responses to crafted probes to determine whether remnants of forgotten data persist. However, such behavioral signatures can be neutralized. Adversaries may train the model adversarially to suppress fingerprint activation (Bai et al., 2021; Jia et al., 2022), or apply input filtering and sanitization mechanisms to block suspicious fingerprint queries (Wang et al., 2022; Abusnaina et al., 2021). Targeted fingerprint removal techniques (Zheng et al., 2022; Guo et al., 2021) can eliminate verification triggers while preserving memorized content, thereby invalidating the behavioral cues relied upon by the auditor.
  2. (2)
    Obfuscating Privacy Leakage Signals. Privacy analysis verification typically rely on specific privacy attacks to extract identifiable privacy signals associated with forgotten data, in order to assess whether such data continues to influence the model. However, these signals can be deliberately obfuscated. For example, retraining with differential privacy (Jia et al., 2019), injecting noise into gradients, or clipping model outputs (Gong et al., 2023; Wen et al., 2021) can all reduce the effectiveness of such analysis. As a result, a model may appear to have forgotten certain data, but this effect may be due to defense mechanisms suppressing the adversary’s ability to identify the forgotten data, rather than actual data removal.
  3. (3)
    Spoofing Performance Degradation. Model performance methods monitor performance metrics to infer whether unlearning has occurred. Yet these indicators can be spoofed by adversaries. Through data poisoning (Tolpegin et al., 2020; Fang et al., 2020), the model’s performance can be manipulated to mimic the expected degradation or improvement patterns. In such cases, the model passes the performance check without performing actual data removal, undermining the verification’s trustworthiness.
@@ 7. Open Questions and Future Directions
A fundamental challenge in unlearning research is the absence of a unified and operational definition of forgetting. Current approaches rely on diverse criteria, including fingerprint (Zhang et al., 2024d; Tam et al., 2024; Goel et al., 2022), privacy analysis (Tu et al., 2024; Di et al., 2025; Sharma et al., 2024), and model performance (Graves et al., 2021; Mehta et al., 2022). This definitional fragmentation creates a critical ambiguity: if a model passes certain verification protocols but fails others, can it still be considered to have truly forgotten the data? Without a common standard, service providers may selectively adopt favorable verification strategies and claim compliance without achieving the intended privacy guarantees. The lack of formal consensus not only impedes fair comparison across methods but also opens the door to unverifiable or misleading claims. Future research should aim to establish a rigorous, interpretable, and widely applicable definition of forgetting that accommodates various verification paradigms and practical deployment needs.
Reliable Unlearning Verification.
Beyond definitional ambiguity, the reliability of existing verification methods remains a pressing concern. Many behavioral approaches, such as those based on privacy analysis or fingerprint, depend heavily on model outputs and the effectiveness of auxiliary attack modules. When these modules are poorly calibrated or overly sensitive to distributional shifts, the resulting verification can become unstable or misleading. At the same time, theoretically grounded methods, including differential privacy verification and reproducibility verification, also suffer from fundamental limitations. For instance, reproducibility verification assumes that if a model trained with and without a specific data point yields identical parameters, then the data has been successfully forgotten. However, as shown in (Thudi et al., 2022), it is possible to construct datasets that differ by a single point $\mathbf{x}$ yet result in nearly indistinguishable models. This allows for false claims of forgetting without any actual data removal. Although recent works (Weng et al., 2024; Eisenhofer et al., 2025) have explored the integration of cryptographic primitives into reproducibility verification to enhance its trustworthiness, these techniques remain computationally intensive and difficult to scale. This underscores the need for verification methods that combine theoretical rigor with practical feasibility. Such methods should provide reliable forgetting guarantees without imposing excessive computational or access requirements.
@@
Despite growing interest in unlearning verification, many methods are developed under overly idealized assumptions. It is often presumed that data distributions are stationary (Chen et al., 2021; Fan et al., 2024), model internals are fully accessible (Guo et al., 2020; Zhang et al., 2024b), and the training process is entirely reproducible (Thudi et al., 2022; Eisenhofer et al., 2025). These assumptions rarely hold in real-world deployments, where models are subject to stochastic training dynamics, evolving input distributions, and limited access due to privacy policies or API restrictions. Such constraints introduce substantial uncertainty that current methods are ill-prepared to handle. For instance, small perturbations in initialization or mini-batch order can significantly affect model outputs, undermining the reliability of behavior signals. Similarly, black-box environments make gradient-based or log-based verification infeasible. This gap between theoretical design and deployment reality remains one of the least addressed yet most impactful challenges. Without explicitly accounting for these constraints, verification results may become unstable, misleading, or fundamentally inapplicable. Future research must incorporate robustness into the core design of verification protocols, including tolerance to randomness, adaptability to shifting distributions, and operability under partial or noisy access. Otherwise, verification will remain a fragile component that functions only in controlled settings and silently fails in practical use.
@@
Another insufficiently explored limitation is the inconsistency of verification effectiveness across different data samples. Many behavioral methods, particularly those based on privacy analysis or perform exceptionally well on atypical or poorly generalized samples, which are the most likely to be memorized by the model (Zhong et al., 2024; Feldman and Zhang, 2020). However, they often underestimate the forgetting failure rate on well-generalized samples, leading to incomplete or biased assessments. This inconsistency reflects a deeper issue: current verification techniques often implicitly assume that all samples are equally verifiable, ignoring the heterogeneous learning dynamics inherent to modern models. Without accounting for sample-specific factors such as gradient influence, model sensitivity, or generalization difficulty, these methods risk producing misleading evaluations of unlearning effectiveness. Future work should aim to develop verification strategies that support consistent and fair assessments across a broad spectrum of data instances.
@@ 8. Conclusion
Machine unlearning is a process that enables machine learning models to forget specific training data through specialized techniques. While these techniques provide algorithm-level forgetting or approximate forgetting, the honesty of the service providers offering unlearning APIs often remains questionable. Therefore, verification becomes a crucial element to enhance the reliability of machine unlearning. However, existing works lack a comprehensive and thorough survey of this field. To fill this gap, we present the first survey on machine unlearning verification. We classify existing verification methods into behavioral and parametric types, analyzing their respective advantages and limitations. Additionally, we examine the associated vulnerabilities and threats and identify several open research questions. We believe that our work can advance the study of the reliability and security of machine unlearning, thus enabling machine learning systems to more reliably handle the forgetting of sensitive data.